If 14.5 grams of sodium nitrate dissolves in 100 cubic centimeters of water and the temperature goes from 26 degrees Celsius to 23 degrees Celsius, calculate : a) the change in temperature (Make sure to write the units just like seen in the problem) *
1
Expert's answer
2020-03-24T14:01:36-0400

Solubility=14.5g/100cc

change in temperature, "dT= Tf-Ti"

Tf=230C

Ti=260C

dT=23-26 =-30C